Decompose



(i)



(x^2+X+1)/(X+3)(x^2-x+1)



into partial fractions (show all the steps).


1
Expert's answer
2022-04-07T17:04:09-0400

"\frac{x^2+x+1}{(x+3)(x^2-x+1)}=" "\frac{A}{x+3}+\frac{Bx+C}{x^2-x+1}"

"x^2+x+1=A(x^2-x+1)+(Bx+C)(x+3)"

Equating coefficients, we get:

"1=A+B" (1)

"1=-A+3B+C" (2)

"1=A+3C" (3)

Substituting (1) into (2) gives

"1=-A+3(1-A)+C=-4A+C+3" (4)

Now let’s substitute (3) into (4)

"1=-4(1-3C)+C+3=13C-1"

"C=\frac{2}{13}"

"A=1-3C=1-\frac{6}{13}=\frac{7}{13}"

"B=1-A=1-\frac{7}{13}=\frac{6}{13}"

"\frac{x^2+x+1}{(x+3)(x^2-x+1)}=""\frac{1}{13}\frac{7}{x+3}+\frac{1}{13}\frac{6x+2}{x^2-x+1}"